# MTVQA: Benchmarking Multilingual Text-Centric Visual Question Answering

### v3

@@ 4.2 Implementation Details
We conduct the evaluation experiments over the baseline MLLMs with their default settings, ignoring the effect of generation configuration on the results. To make the output of MLLMs more evaluation-friendly, we design the following prompt format to limit the output length: “Answer the question using a word or phrase in the language of the question. + <Question>”, where “<Question>” represents the actual question from the MTVQA test set. This approach aims to make the answers as concise as possible. Besides, InternLM-Xcomposer2-4KHD  [11] is chosen as the base model for an instruction tuning experiment on the MTVQA training set. The instruction tuning process adheres to the default training settings specified by the source, with “HD-16” and completes one epoch of training on 8 NVIDIA-A100 GPUs within 2 hours.
@@
Table 3: Few-shot performance of GPT-4V on the MTVQA benchmark. In-context examples are randomly selected from the training set of MTVQA in the respective languages. “n-shot” represents the number of the selected in-context examples.
@@
Table 4: Comparison on applying OCR to GPT-4 and GPT-4V
AR DE FR IT JA KO RU TH VI Avg.
OCR+GPT-4 18.9 21.2 29.8 27.3 15.8 24 10.6 19 28.2 21.6
OCR+GPT-4V 22.3 35.1 42.6 36.7 19.2 32.1 12.1 20.9 34.1 28.3
GPT-4V 11.5 31.5 40.4 32.3 11.5 16.7 10.3 15 28.9 22.0
4.3 Evaluation Results
Evaluation metric. To accurately assess whether the visual text that occurs in the answer is correct, we adopt Accuracy as the metric. The Accuracy metric measures the percentage of questions for which the predicted answer matches with any of the target answers for the question. We also compare the models and human performance using the ANLS and Accuracy metrics. The models perform a little higher (5% - 10%) when using ANLS, but the rankings are mostly unchanged. Since the ANLS metrics do not accurately reflect whether the correct textual content in the image is being presented, we chose the Accuracy metric.
Human performance. To establish a benchmark for human performance on MTVQA, 10 well-educated native speakers are randomly selected for each language. As shown in Table 2, the average human performance on MTVQA is around 80%.
[…]
@@ 4.3 Evaluation Results
Few-shot evaluation of GPT-4V. Here, we compare the performance of GPT-4V on MTVQA under the few-shot settings. Specifically, we perform zero-shot, two-shot, five-shot, and eight-shot evaluations. We randomly select in-context examples from the train set in the respective languages and evaluate GPT-4V on the remaining instances. As shown in Tab. 3, compared to the zero-shot setting, GPT-4V’s performance has improved considerably under few-shot, highlighting its exceptional in-context learning ability in multilingual text comprehension contexts. Moreover, comparisons based on varying numbers of in-context samples reveal that augmenting the in-context samples can aid in further enhancement; however, after reaching a certain volume, the improvement becomes saturated.
Experiments on Text-Based LLMs and MLLMs with OCR results. To compare text-based LLMs and MLLMs performance on the MTVQA bench with OCR results, we adopt the Bytedance multilingual OCR API tool to extract multilingual text from images. For text-based LLMs, we choose GPT-4, and for MLLMs, we choose GPT-4V. As shown in Table  4, GPT-4V and OCR+GPT-4 have similar performance, but both are much lower than OCR+GPT-4V. We analyze the cases in detail and find differences in the challenges encountered by OCR+GPT-4 and GPT-4V. Since the performance of Latin languages is weaker than that of GPT4V(strong Latin text perception and comprehension abilities), OCR+GPT4 issues lie more in the lack of perception of visual elements and position. Many questions are dependent on visual elements and positional relationships in the image, resulting in that cannot be answered. GPT-4V issues lie more in the weakness of perception and comprehension of visual text, especially non-Latin text (AR/JA/KO/TH in Table 4), but GPT-4V can capture visual elements and positional relationships in images. More importantly, all three settings perform poorly and the comprehension of multilingual visual text remains a challenging case.
5 Limitation
The current version of the MTVQA dataset, while dialectically diverse, exhibits limitations in its language coverage. Despite encompassing a range of languages, it falls short of inclusivity, omitting numerous lesser-spoken languages. This prompts our future continual endeavor to ensure comprehensive representation across the linguistic spectrum.
6 Conclusion
To attack the visual-textual misalignment issue in multilingual TEC-VQA, we introduce MTVQA, a new benchmark featuring high-quality human expert annotations in 9 diverse languages. We believe that MTVQA is the first benchmark to provide fully manual annotations tailored to text-centric scenarios. The results obtained from closed-source, general-purpose, and text-centric MLLMs on our MTVQA dataset indicate that there is still room for improving their performance in multilingual text-centric scenarios. Although the MTVQA dataset has limitations, such as the underrepresentation of several lesser-spoken languages and single canonical answers, future updates will address these issues by expanding the multilingual scope and including a range of plausible answers. We are confident that this dataset can inspire researchers within the TEC-VQA community with new perspectives and ideas.

### v4

@@ 4.2 Implementation Details
We conduct the evaluation experiments over the baseline MLLMs with their default settings, ignoring the effect of generation configuration on the results. To make the output of MLLMs more evaluation-friendly, we design the following prompt format to limit the output length: “Answer the question using a word or phrase in the language of the question. + <Question>”, where “<Question>” represents the actual question from the MTVQA test set. This approach aims to make the answers as concise as possible. Besides, InternLM-Xcomposer2-4KHD  (Dong et al., 2024) is chosen as the base model for an instruction tuning experiment on the MTVQA training set. The instruction tuning process adheres to the default training settings specified by the source, with “HD-16” and completes one epoch of training on 8 NVIDIA-A100 GPUs within 2 hours.
models GPT-4V GPT-4o Gemini Claude3 Qwen2-VL 72B InternVL2.5 78B
ANLS 29.6 36.9 31.7 35.4 38.7 41.6
Acc 22.0 27.8 23.2 25.7 30.9 32.2
Table 3: Model performance comparison under ANLS and Acc metric.
AR DE FR IT JA KO RU TH VI Avg.
OCR+GPT-4 18.9 21.2 29.8 27.3 15.8 24 10.6 19 28.2 21.6
OCR+GPT-4V 22.3 35.1 42.6 36.7 19.2 32.1 12.1 20.9 34.1 28.3
GPT-4V 11.5 31.5 40.4 32.3 11.5 16.7 10.3 15 28.9 22.0
Table 4: Comparison on applying OCR to GPT-4 and GPT-4V.
4.3 Evaluation Results
4.3.1 Evaluation metric
To accurately assess whether the visual text that occurs in the answer is correct, we adopt Accuracy as the metric. The Accuracy metric measures the percentage of questions for which the predicted answer matches any of the target answers for the question. We also compare the models and human performance using the ANLS and Accuracy metrics. The models perform a little higher (5% - 10%) when using ANLS(Tab.3), but the rankings are unchanged. Since the ANLS metrics do not accurately reflect whether the correct textual content in the image is being presented, we chose the Accuracy metric.
4.3.2 Human performance
To establish a benchmark for human performance on MTVQA, 10 well-educated native speakers are randomly selected for each language. As shown in Table 2, the average human performance on MTVQA is around 80%, which is far beyond the performance of state-of-the-art MLLMs. It demonstrates the shortcomings and the huge room for improvement of the existing models in multilingual visual text perception and comprehension.
4.3.3 Zero-shot evaluation
[…]
4.3.4 Instruction tuning
[…]
4.3.5 Experiments on Text-Based LLMs and MLLMs with OCR results
To compare text-based LLMs and MLLMs’ performance on the MTVQA bench with OCR results, we adopt the Bytedance multilingual OCR API tool to extract multilingual text from images. For text-based LLMs, we choose GPT-4, and for MLLMs, we choose GPT-4V. As shown in Table  4, GPT-4V and OCR+GPT-4 have similar performance, but both are much lower than OCR+GPT-4V. We analyze the cases in detail and find differences in the challenges encountered by OCR+GPT-4 and GPT-4V. Since the performance of Latin languages is weaker than that of GPT4V(strong Latin text perception and comprehension abilities), OCR+GPT4 issues lie more in the lack of perception of visual elements and position. Many questions are dependent on visual elements and positional relationships in the image, resulting in which cannot be answered. GPT-4V issues lie more in the weakness of perception and comprehension of visual text, especially non-Latin text (AR/JA/KO/TH in Table 4), but GPT-4V can capture visual elements and positional relationships in images. More importantly, all three settings perform poorly, and the comprehension of multilingual visual text remains a challenging case.
4.3.6 Failure Case Analysis
We have compiled preliminary statistics on common errors: OCR failures (39%), reasoning gaps (34%), language bias (15%), and hallucinations (12%). For example, OCR Failure: Model misreads "café" as "cafe" in a French menu. Reasoning Gap: Fails to infer "9:00–16:00" as "business hours" in a Russian poster.
5 Dataset Licenses
The MTVQA dataset could be used publicly under the CC BY-NC 4.0 license.
6 Acknowledgement
This work is supported by the Natural Science Foundation of China (NSFC) NO. 62225603.
7 Conclusion
To attack the visual-textual misalignment issue in multilingual TEC-VQA, we introduce MTVQA, a new benchmark featuring high-quality human expert annotations in 9 diverse languages. We believe that MTVQA is the first benchmark to provide fully manual annotations tailored to text-centric scenarios. The results obtained from closed-source, general-purpose, and text-centric MLLMs on our MTVQA dataset indicate that there is still room for improving their performance in multilingual text-centric scenarios. Although the MTVQA dataset has limitations, such as the underrepresentation of several lesser-spoken languages, future updates will address these issues by expanding the multilingual scope and including a range of plausible answers. We are confident that this dataset can inspire researchers within the TEC-VQA community with new perspectives and ideas.
8 Limitation
The current version of the MTVQA dataset, while dialectically diverse, exhibits limitations in its language coverage. Despite encompassing a range of languages, it falls short of inclusivity, omitting numerous lesser-spoken languages. This prompts our future endeavor to ensure comprehensive representation across the linguistic spectrum.
Appendix A More Experiments
@@
Table 5: Few-shot performance of GPT-4V on the MTVQA benchmark. In-context examples are randomly selected from the training set of MTVQA in the respective languages. “n-shot” represents the number of selected in-context examples.
@@
A.1 Few-shot evaluation of GPT-4V.
@@
Here, we compare the performance of GPT-4V on MTVQA under the few-shot settings. Specifically, we perform zero-shot, two-shot, five-shot, and eight-shot evaluations. We randomly select in-context examples from the training set in the respective languages and evaluate GPT-4V on the remaining instances. As shown in Tab. 5, compared to the zero-shot setting, GPT-4V’s performance has improved considerably under few-shot, highlighting its exceptional in-context learning ability in multilingual text comprehension contexts. Moreover, comparisons based on varying numbers of in-context samples reveal that augmenting the in-context samples can aid in further enhancement; however, after reaching a certain volume, the improvement becomes saturated.
A.2 MLLM Performance Changes When the Question is Asked in English.
We translate the questions into English using GPT-4 and then perform a test on GPT4V. The result is shown in Table  6. The GPT4V performance is robust when the question is asked in English and the original language.
Table 6: GPT4V performance changes when the question is asked in English
AR DE FR IT JA KO RU TH VI Average
English Questions 11.9 31.8 39.0 32.1 12.0 15.8 10.1 18.1 28.5 22.1
Original Questions 11.5 31.5 40.4 32.3 11.5 16.7 10.3 15.0 28.9 22.0
A.3 The Exploration of Additional Metrics
We compare ANLS (soft metric) and Accuracy, and we choose Accuracy to reflect the consistency of the answer and the visual text content because accurate answers (e.g., price, date) are critical. We also pilot LAVE but observe low consistency (76.8%) with human judgments for non-Latin languages (Table 2), due to the lower low-resource linguistic performance of LLMs.
Appendix B More Visualizations
(a) German (DE) example
(b) Italian (IT) example
(c) French (FR) example
(d) Japanese (JA) example
(e) Korean (KO) example
Figure 6: Multilingual VQA examples selected from five languages. The corresponding English translations are in brackets.
Table 7: Mean lengths of question-answer pairs in different languages of the training set and test set, using GPT-4o tokenizer.
AR DE FR IT JA KO RU TH VI
Training Set
Question 8.29 8.72 9.73 12.05 12.43 11.74 11.56 11.35 11.21
Answer 9.66 6.96 7.34 11.24 12.70 13.56 12.00 11.26 13.31
Test Set
Question 8.08 8.29 9.76 11.93 12.48 12.2 11.65 10.98 10.99
Answer 7.95 6.67 6.61 11.04 12.55 13.61 14.42 12.08 13.00
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 7: Word clouds showcasing top answers in various languages, tokenized via NLTK with removing stop words, punctuation, and digits.